`please help me solve each system of equation
1.) x-2y=-14 2.) x+2y=4 3.) 3x-y=2
y+2x=4 2x+y=-1 y=2x-5
and whats the solution??'
Can you help me solve all of them is because i am bad at math ={ Thank You I would really appreciate it

Answer by checkley77(12844)

1)x-2y=-14 OR x=2y-14 now replace x with(2y-14) in the other equation & solve for y.
y+2x=4
y+2(2y-14)=4
y+4y-28=4
5y=4+28
5y=32
y=32/5 ans
x=2*32/5-14
x=64/5-14
x=(64-14*5)/5
x=(64-70)/5
x=-6/5 ans.
Proof:
32/5+2*-6/5=4
32/5-12/5=4
20/5=4
4=4
2)x+2y=4 or x=4-2y now replace x with (4-2y) in the other equation & solve for y.
2x+y=-1
2(4-2y)+y=-1
8-4y+y=-1
8-3y=-1
-3y=-1-8
-3y=-9
y=-9/-3
y=3 ans.
x=4-2*3
x=4-6
x=-2
Proof:
-2+2*3=4
-2+6=4
4=4
3)3x-y=2
y=2x-5 replace y with (2x-5) in the other equation & solve for x.
3x-(2x-5)=2
3x-2x+5=2
x=2-5
x=-3 ans.
y=2*-3-5
y=-6-5
y=-11 ans.
Proof:
3*-3-(-11)=2
-9+11=2
2=2